Growth​ Company's current share price is $20.00 and it is expected to pay a $1.25 dividend per share next year. After​ that, the​ firm's dividends are expected to grow at a rate of 4.2% per year.

a. What is an estimate of Growth​ Company's cost of​ equity?

b. Growth Company also has preferred stock outstanding that pays a $2.25 per share fixed dividend. If this stock is currently priced at $28.15​, what is Growth​ Company's cost of preferred​ stock?

c. Growth Company has existing debt issued three years ago with a coupon rate of 5.5%. The firm just issued new debt at par with a coupon rate of 6.1%. What is Growth​ Company's cost of​ debt?

d. Growth Company has 5.1 million common shares outstanding and 1.2 million preferred shares​ outstanding, and its equity has a total book value of $49.9 million. Its liabilities have a market value of $19.5 million. If Growth​ Company's common and preferred shares are priced as in parts (a​) and ​(b​), what is the market value of Growth​ Company's assets?

e. Growth Company faces a 40% tax rate. Given the information in parts ​(a​) through ​(d​), and your answers to those​ problems, what is Growth​ Company's WACC?

​Note: Assume that the firm will always be able to utilize its full interest tax shield.
Answer #1

a. Cost of equity =Dividend Next Year/Price +Growth =1.25/20+4.2% =10.45%
b. Cost of Preferred Dividend =Dividend/Price =2.25/28.15 =7.99%
c.Since debt is sold at par , Hence cost of debt =6.1%
d, Market value of common shares =Number of Shares*Share Price =5100000*20 =102,000,000 or 102 million
Market value of Preferred Sock =Number of Shares*Share Price of Preferred Stock =1200000*28.15 =33780000 or 33.78 million
Market Value of Liabilities =19.5 million
Market Value of Total Assets =102+33.78+19.5 =155.28 million